OpenStudy (anonymous):

can anyone integarte and simplify 2x^3-3x/4x

OpenStudy (kropot72):

Is this the integral to be solved? \[\int\limits_{}^{}((2x ^{3}-3x)\div(4x))dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Assuming you meant \[ \frac{ 2x^3 - 3x}{4x} = \frac{2x^3}{4x} - \frac{3x}{4x} \], where we can integrate the two bit : 2x^3/4x has for integral x^3/6 and 3x/4x has 3x/4 (respectively from the polinomial+quotient and quotient rule ). So the final result is -((3 x)/4) + x^3/6 which can be factored as (1/12) (x) (-9 + 2 x^2)

OpenStudy (campbell_st):

simplify 1st \[\int\limits 2x^3/4x - 3x/4x dx = \int\limits x^2/2 - 3/4 dx\]

OpenStudy (kropot72):

If both terms in the denominator are divided by 4x the result is: \[\int\limits_{}^{}((x ^{2}/2)-(3/4))dx=(x ^{3}/6)-(3x/4)+C\]

OpenStudy (campbell_st):

this becomes \[1/2 \times x^3/3 3x/4 + C = x^3/6 - 3x/4 + C\]

OpenStudy (kropot72):

My reply should have said"If both terms in the numerator.............". Sorry for the slip:)